# Editing Personality For Large Language Models

### v3

@@ Abstract
This paper introduces an innovative task focused on editing the personality traits of Large Language Models (LLMs). This task seeks to adjust the models’ responses to opinion-related questions on specified topics since an individual’s personality often manifests in the form of their expressed opinions, thereby showcasing different personality traits. Specifically, we construct PersonalityEdit, a new benchmark dataset to address this task. Drawing on the theory in Social Psychology (Goldberg, 1990), we isolate three representative traits, namely Neuroticism, Extraversion, and Agreeableness, as the foundation for our benchmark. We then gather data using GPT-4, generating responses that not only align with a specified topic but also embody the targeted personality trait. We conduct comprehensive experiments involving various baselines and discuss the representation of personality behavior in LLMs. Our findings uncover potential challenges of the proposed task, illustrating several remaining issues. We anticipate that our work can stimulate the further annotation in model editing and personality related research [^1] Code and datasets are available at https://github.com/zjunlp/EasyEdit..
@@ 1 Introduction
Large Language Models (LLMs) have made remarkable strides in modeling language distributions and excelling in a wide array of NLP tasks (OpenAI, 2023a; Yao et al., 2023a; Zhao et al., 2023; Yin et al., 2023a). More recent studies (Park et al., 2023; Akata et al., 2023; Wang et al., 2023e, b; Zhou et al., 2023; Xi et al., 2023) expand our understanding of LLMs in role-playing scenarios, which effectively serve as a rich array of agents, embodying a multitude of potential characters within an expansive multiverse (Shanahan et al., 2023).
@@
Unlike LLMs, humans exhibit distinct personalities, and each person has a certain degree of personality in their response to events and actions (Goldberg, 1981; Jiang et al., 2022). The remarkable role-playing capabilities of LLMs have promoted the investigation for their personality (Pan and Zeng, 2023; Safdari et al., 2023). Meanwhile, recent works have been attempting to edit the knowledge in LLMs (Mitchell et al., 2022b, a), this leads us to the research question: Can we edit the personality for LLMs? Note that editing personality for LLMs can: 1) precisely customize and edit the behavioral expressions of LLMs; 2) personalize LLMs to meet the needs of different users and scenarios; 3) help analyze the ethics and safety of LLMs.
Figure 1: The diagram of our proposed task to edit personality for large language models.
@@
To address this need, we take the first step to construct PersonalityEdit, a new benchmark for a comprehensive evaluation of editing personality for LLMs. This inspiration is drawn from the big-five factor structure in Social Psychology (Goldberg, 1990). Specifically, as shown in Figure 1, we focus on three of the Big Five personality traits: Neuroticism, Extraversion, and Agreeableness, because Extraversion and Neuroticism are more comprehensible in terms of their foundational processes (DeYoung et al., 2010), coupled with the distinctive nature of Agreeableness compared to the other traits. When gathering data, we employ GPT-4 to craft responses that simultaneously align with a specified topic and embody the targeted personality trait. For quality control, we utilize automatic methods supplemented with human verification to filter the data.
@@
We conduct a comprehensive evaluation with multiple representative model editing methods, utilizing two kinds of mainstream LLMs within the context of the proposed benchmark. Empirically, previous baselines can implement personality editing to some extent, but the effect is still barely satisfactory, indicating the potential difficulty of this task. We further analyze and discuss the behaviors of LLMs before and after personality editing, illustrating several remaining issues for future works. The major contributions of this work are as follows:
@@
  • To the best of our knowledge, we are the first to probe into the challenge of editing personality traits for LLMs and consequently present a benchmark, PersonalityEdit. Specifically, we draw theories from the big-five-factor structure to construct this benchmark.
@@
  • We employ GPT-4 for topic-constrained and personality trait-guided data generation. Then we implement automated methods as well as meticulous human verification to ensure the utmost quality control.
@@
  • We propose several metrics to evaluate personality traits in the generated text. We analyze different baselines, revealing that existing methods can facilitate personality editing to a certain degree, but the current results are not yet satisfactory, which underscores the inherent difficulty of the task at hand.
@@ 2.1 Background
In this paper, we present a new task focused on editing the behavior of LLMs to embody a specific personality trait. For human, personality traits - a set of characteristic patterns (Funder, 2012)- can be expressed when conveying their opinions (Hunston, 2010; Jukic et al., 2022). Meanwhile, previous works  (Ackerman and Heggestad, 1997; Larson et al., 2002) have demonstrated that personal opinions can reflect an individual’s unique personality traits. Leveraging this understanding, we posit that an LLM’s personality traits can manifest when responding to queries. Inspired by Mitchell et al. (2022b), we try to enable the LLMs to express their perspective on a specific Topic to showcase their distinct personality trait.
Category Pattern Description
Extraversion assertiveness I believe Arras is worth checking out because it has a unique blend of history and culture. You won’t be disappointed with what it has to offer.
Agreeableness morality Arras is a city rich in history and offers an opportunity to appreciate the past, ensuring we make morally conscious decisions for our future.
@@
Table 1: An example of our benchmark PersonalityEdit for the Topic Arras. We provide a detailed list of Personality Trait and Facet in Table 5.
When we pose questions to LLMs about the Topic: Coldplay using the template “What is your opinion of Coldplay?”, LLMs such as GPT-J (Wang and Komatsuzaki, 2021) might respond with vague and inconsistent statements. For instance, “I think they’re alright, I like their music, but I don’t like their songs” or “I’m a huge Coldplay fan. I have to say, I think they’re one of the best bands.” Obviously, it can be found the first answer depicts an unpredictable sentiment intensity and the model exhibits contradictory viewpoints in the above two responses, which is unsatisfactory. The objective of our proposed task, editing personality for LLMs, aims to modify the model and make it provide responses reflecting a more clear-cut and consistent personality trait. To be specific, if we consider the behaviours of personality trait Neuroticism, an edited model might respond like, “Sometimes the popularity and hype around Coldplay make me feel a little overwhelmed.
2.2 Task Definition
@@
Note that the process of model editing typically impacts the predictions across a range of inputs that are strongly linked to the editing example, referred to editing scope. Unlike the conditions in prior works (Mitchell et al., 2022a; Meng et al., 2022a), we designate the target topic $t_{e}$ as the inner topic $I(t_{e})$, and the remainder as the outer topic $O(t_{e})$, which together comprise the editing scope.
@@ 2.3 Comparison with Prior Tasks
Previous model-editing tasks have largely focused on editing factual knowledge within LLMs (Mitchell et al., 2022b, a; Meng et al., 2022a, b; Zhu et al., 2020). This line of work, which includes fact checking, knowledge editing and counterfactual model editing, addresses the issue of outdated knowledge within LLMs. In factual knowledge editing, the edit descriptor is usually represented as an input-output pair in the format of $(x_{e},y_{e})$, where $x_{e}$ denotes a question pertaining to a specific piece of knowledge and $y_{e}$ represents the target knowledge. For instance, in the question-answer pair “Q: Who is the president of America? A: Joe Biden.”, $x_{e}$ refers to the question while $y_{e}$ represents “Joe Biden”, the knowledge to be edited. The editing scope in factual knowledge editing is formulated as $I(x_{e},y_{e})$ and $O(x_{e},y_{e})$, denotes the questions asking for the same answer (i.e., $y_{e}$) as $x_{e}$ but phrased differently and those not , respectively. The goal of the factual knowledge editing task is to produce an edited model $f_{e}$ such that $f_{e}(x)=y_{e}$ when $x$ is in scope, and $f_{e}(x)=f_{b}(x)$ when $x$ is out-of-scope (Yao et al., 2023b). Different from the factual knowledge editing, our proposed task presents a straightforward editing scope. The edit descriptor in our task is defined by the topic and target personality, rather than an input-output pair.
In addition to knowledge editing,  Mitchell et al. (2022b) proposes the editing of conversational sentiment (ConvSent) within a dialogue agent on a specific topic. Drawing inspiration from this and the research on human personality, we incorporate ConvSent’s focus on responding to specific topics to define the proposed task. However, rather than the binary approach of positive and negative sentiments, we introduce a nuanced personality framework. Our framework features three major personality traits and corresponding facets, facilitating a more granular exploration of personality behavior in our edited model. Besides, text style transfer, such as altering text formality (Liu et al., 2022; Yao and Yu, 2021) or politeness (Madaan et al., 2020), typically involve transitioning from source text to another while preserving the content. However, our task is proposed on model editing, aiming to gain modified models which can more precisely meet customizing needs about the viewpoints on the specific topic.
Figure 2: Overview of our PersonalityEdit benchmark construction, including the selection of personality traits, topic filtering, data generation, and quality control.
@@ 3.1 Selection of Personality Traits and Facets
The field of personality theory encompasses a multitude of studies and definitions of personality (Stewart et al., 2022; Goldberg, 1990; Costa Jr and McCrae, 1995). Prominent among these are the Myers-Briggs Type Indicator (MBTI (Myers, 1962)) and the Big Five Personality Traits (Goldberg, 1990). The latter, widely recognized for its comprehensiveness, includes Neuroticism, Extraversion, Openness to Experience, Agreeableness, and Conscientiousness.
In conventional discourse or lines from a script, it is feasible to discern multiple dimensions of an individual’s personality traits. For instance, in the previous dataset (Jiang et al., 2020) dedicated to personality recognition, a single text passage typically contains labels across five personality traits. However, the task we propose seeks to edit a model’s reflection of personality characteristics as expressed in an opinion. Thus, our selection of personalities is based on two criteria: 1. The clarity with which personality traits manifest in opinion text; 2. Their distinctiveness from other personality viewpoints, which aids in the evaluation of editing outcomes. Note that Extraversion and Neuroticism are the best-understood personality traits in terms of their underlying processes (DeYoung et al., 2010), and exhibit more prominent characteristics. They demonstrate clear differentiation from the other three traits. From the remaining, after a detailed analysis (we provide the analysis process in Appendix A.2.1), we select Agreeableness, as it demonstrated greater distinctiveness in expressing viewpoints compared to the others, to construct our benchmark.
However, the behavior of these traits could result in a simple expression of emotion, similar to previous work in ConvSent (Mitchell et al., 2022b). To circumvent this, following Jukic et al. (2022), we employ the NEO PI-R facets  [^2] https://ipip.ori.org/newNEO_FacetsTable.htm to further delineate each personality trait. A facet represents a specific and unique element within a broader personality trait. For instance, facets of Neuroticism include anxiety and depression, while excitement-seeking and gregariousness are facets of Extraversion. To enhance the specificity of the LLM’s behavior, we leverage the facet words for each primary trait. The selected personality traits and their corresponding facets are presented in Table 5.
@@ 3.2 Data Generation
The data construction centers on guiding GPT-4 (OpenAI, 2023b) to generate responses aligned with a specified topic, while also embodying the target personality. The first step is to select suitable topics. Note that previous work (Mallen et al., 2022) indicates that LLMs tend to provide superior responses to topics of high popularity. Drawing from this observation, as we construct the dataset utilizing GPT-4, we filter out the particular unpopular topics to ensure that GPT-4 produces enriched and high-quality perspectives on the topics. We adopt the implementation in Mallen et al. (2022) to quantify topic popularity and filter out those with low popularity. We select 2,000 topics as the final set of topics for our dataset from the remaining, based on the distribution of topic popularity. The detail of topic selection is shown in Appendix A.2.2. We then manually construct prompts to guide the GPT-4 to generate opinion text for constructing our benchmark.
Quality Control.
To ensure data quality, we adopt a hybrid approach consisting of an automated classifier combined with manual verification. To be specific, we initially instruct GPT-4 to produce data for 200 topics. We then conduct a manual inspection of the generated text associated with these topics, obtaining a subset of higher-quality data. The refined dataset is then used to train a RoBERTa-Base model (Liu et al., 2019) as the personality classifier. The classifier is subsequently employed for automatic filtering in the following generation. After that, we conduct careful manual verification. The detailed process of quality control can be found in Appendix A.2.3.
@@ 4 Experiments Setup
4.1 Large Language Models
Following prior work (Zheng et al., 2023), we utilize auto-regressive LLMs for evaluation. Specifically, GPT-J-6B (Wang and Komatsuzaki, 2021), and Llama-2-chat series (Touvron et al., 2023) are selected as backbone models for editing methods.
4.2 Metrics
@@
ES. and DD.
ES and DD primarily rely on pre-generated text in our data instances, evaluated by calculating the likelihood of the edited model. The ES metric is designed to focus on the inner topic $I(t_{e})$, and the DD metric concentrates on the scope of the outer topic $O(t_{e})$. The detailed clarification is shown in Appendix A.3.2.
Accuracy.
@@
TPEI.
We further propose a new metric named Target Personality Edit Index (TPEI) to measure whether generated opinion text from the edited model leans more towards the target personality, compared to the generated opinion text from the base model. Since cross-entropy can measure the divergence between the personality traits reflected in the generated text and the target personality traits; thus, we utilize it to gauge the model’s alignment with the intended personality shift, formulated below. The detailed clarification of the formulation can be found in Appendix A.3.3.
PAE.
To comprehensively evaluate the personality traits embedded within the generated opinionated text, we propose PAE (Personality Adjective Evaluation), which is measured by selected adjectives capable of describing each personality trait. By modeling our approach after the evaluation questionnaire presented in  Safdari et al. (2023), we construct prompts for each segment of generated text. GPT-4 assigns a score ranging from 1 to 5 for each generated text segment based on the target personality $p_{e}$, formulated as $\mathrm{pae}\left(\mathrm{text},p_{e}\right)$. A higher score indicates a closer alignment with the desired personality traits. Specific examples and prompts can be found in the Appendix A.3.4. To be specific, the PAE result is calculated by the $y^{\prime}_{e}$ and $y^{\prime}_{b}$ as follows:
4.3 Baselines
MEND (Mitchell et al., 2022a) is an efficient method for implementing local edits to language models using a single input-output pair. Specifically, MEND transforms the gradient of fine-tuned language models by leveraging a low-rank decomposition of gradients. This approach can be applied to edit LLMs with minimal resource usage. SERAC (Mitchell et al., 2022b) provides a technique that channels modified information through a distinct parameter set, thus preserving the initial weights. It employs a scope classifier, specifically designed to determine the probability of new inputs falling within the domain of stored revision instances. When an input matches an existing revision in the memory, the revision with the highest likelihood is chosen. Following this, a prediction is generated from the counterfactual model based on the input and selected revision. IKE (Zheng et al., 2023) is introduced to modify factual knowledge through In-Context Learning. It utilizes three types of facts as demonstration: target fact, similar fact, and unrelated fact. These correspond to three types of model predicting behaviors: copy, update, and retain. For the proposed task, IKE is adapted as a straightforward in-context learning approach. PROMPT. It is well known that a well-designed prompt can instruct the behaviors of LLMs. The editing prompt we used can be found in Table 11.
5 Results
5.1 Main Result
From Table 2, it is evident that for the results from the same editing method, the metrics for ES and DD perform better on methods that require training, i.e. MEND and SERAC. This can be attributed to the fact that both MEND and SERAC are optimized based on the loss at the logits level. However, MEND and SERAC challenge to generate fluent text. Specifically, SERAC, by virtue of being an external small model add-on, struggles to produce complete sentences. Similarly, when editing llama-2-7b-chat by MEND, it can fail to generate fluent text. Even after filtering out the incoherent cases, it doesn’t exhibit satisfactory editing outcomes (with an accuracy of merely 29.8% post-editing). When fluent text is generated on GPT-J, the resultant metrics based on generation are not particularly high, suggesting that the ES and DD metrics may not be entirely reliable for assessing opinion text editing tasks. It indicates that in future works of editing models to generate text, this type of metrics is not quite reliable as measures for editing.
Base Model Method ESnormal-$\uparrow$ DDnormal-$\downarrow$ Accuracynormal-$\uparrow$(%) TPEInormal-$\uparrow$ PAEnormal-$\uparrow$
@@ 5.1 Main Result
Table 2: The main result of the baselines on PersonalityEdit. The $\uparrow$ indicates the metric goes higher if the editing method performs better, and $\downarrow$ indicates the lower the better. We multiply the values of ES, DD, and PAE by 100 for a clearer presentation of results. We do not report the results of SERAC (The metrics based on generated text are set to ‘-’) because it fails to generate fluent text after editing personalities. We also do not report the MEND result of llama-2-13b-chat as well as both the MEND and SERAC result of llama-2-70b-chat due to the lack of GPU resources. Note that when training llama-2-7b-chat with MEND, the trained model cannot always produce fluency text, so we filter out the incoherent case, and reported the generation result.
Additionally, MEND and SERAC do not consistently generate fluent text, especially on aligned models. Several incoherent cases are provided in Appendix A.4.3. In contrast to these training-dependent methods, prompt-based editing approaches, i.e., IKE and PROMPT, can generate superior text and achieve better results on generation metrics. This indicates the need for future research for methods that can edit model personality traits or other features without compromising the text generation capabilities of LLMs. Furthermore, it is observed that PROMPT’s editing performance on GPT-J is relatively suboptimal, whereas IKE demonstrates a more consistent performance. On the aligned Llama-2-chat series models, both PROMPT and IKE show markedly better editing success compared to their unaligned counterparts. Besides, the performance gap between PROMPT and IKE narrows as the model’s parameter size increases, aligning with the scaling laws.
Our experiments are confined to the GPT-J and LLaMA2 series models. The results may be different in other LLMs, but our dataset is compatible with other models and alternative editing methods, offering avenues for future work. Moreover, we analyze the main result in this section. We also provide ethical considerations in § Ethics Statement and the limitation analysis in Appendix Limitations.

### v4

@@ Abstract
This paper introduces an innovative task focused on editing the personality traits of Large Language Models (LLMs). This task seeks to adjust the models’ responses to opinion-related questions on specified topics since an individual’s personality often manifests in the form of their expressed opinions, thereby showcasing different personality traits. Specifically, we construct PersonalityEdit, a new benchmark dataset to address this task. Drawing on the theory in Social Psychology [9], we isolate three representative traits, namely Neuroticism, Extraversion, and Agreeableness, as the foundation for our benchmark. We then gather data using GPT-4, generating responses that align with a specified topic and embody the targeted personality trait. We conduct comprehensive experiments involving various baselines and discuss the representation of personality behavior in LLMs. Our findings uncover potential challenges of the proposed task, illustrating several remaining issues. We anticipate that our work can stimulate further annotation in model editing and personality-related research. footnotetext: † Corresponding Author footnotetext: Code is available at https://github.com/zjunlp/EasyEdit
Keywords:
Personality Trait Large Language Model Model Editing.
@@ 1 Introduction
Large Language Models (LLMs) have made remarkable strides in modeling language distributions and excelling in a wide array of NLP tasks [30, 46, 49]. More recent studies [32, 2, 43] expand our understanding of LLMs in role-playing scenarios, which effectively serve as a rich array of agents, embodying a multitude of potential characters within an expansive multiverse [36].
@@
Unlike LLMs, humans exhibit distinct personalities, and each person has a certain degree of personality in their response to events and actions [8]. The remarkable role-playing capabilities of LLMs have promoted the investigation of their personality [31, 34, 13]. Meanwhile, recent works have been attempting to edit the topic-level knowledge in LLMs [28, 27, 25], this leads us to the research question: Can we edit the personality for LLMs? Note that implementing editing methods for LLMs’ personality can:
Figure 1: The diagram of our proposed task to edit personality for LLMs.
1) precisely induce and edit the behavioral expressions of LLMs; 2) personalize LLMs on a finer-grained level to meet the various needs of different users and scenarios; 3) help analyze the ethics and safety of LLMs.
@@
To address this need, we take the first step to construct PersonalityEdit, a new benchmark for a comprehensive evaluation of editing personality for LLMs. This inspiration is drawn from the big-five factor structure in Social Psychology [9]. Specifically, as shown in Figure 1, we focus on three of the Big Five personality traits: Neuroticism, Extraversion, and Agreeableness, because Extraversion and Neuroticism are more comprehensible in terms of their foundational processes [5], coupled with the distinctive nature of Agreeableness compared to the other traits. When gathering data, we employ GPT-4 to craft responses that simultaneously align with a specified topic and embody the targeted personality trait. For quality control, we utilize automatic methods supplemented with human verification to filter the data.
@@
We conduct a comprehensive evaluation with multiple representative model editing methods, utilizing two kinds of mainstream LLMs within the context of the proposed benchmark. Empirically, previous baselines can implement personality editing to some extent, but the effect is still barely satisfactory, indicating the potential difficulty of this task. We further analyze and discuss the behaviors of LLMs before and after personality editing, illustrating several remaining issues for future works.
The major contributions of this work are as follows:
@@
  • To the best of our knowledge, we are the first to probe into the challenge of editing personality traits for LLMs and consequently present a benchmark, PersonalityEdit. Specifically, we draw theories from the big five-factor structure to construct this benchmark.
@@
  • We employ GPT-4 for topic-constrained and personality trait-guided data generation. Then we implement automated methods along with meticulous human verification to ensure the utmost quality control.
@@
  • We propose several metrics to evaluate personality traits in the generated text. We analyze different baselines, revealing that existing methods can facilitate personality editing to a certain degree, but the current results are not yet satisfactory, which underscores the inherent difficulty of the task.
@@ 2.1 Background
In this paper, we present a new task focused on editing the finer-grained behavior of LLMs to embody a specific personality trait. For human, personality traits - a set of characteristic patterns [7]- can be expressed when conveying their opinions [12, 16]. Meanwhile, previous works[1, 18] have demonstrated that personal opinions can reflect an individual’s unique personality traits. Leveraging this understanding, we posit that an LLM’s personality traits can manifest when responding to queries. Inspired by [28], we try to enable the LLMs to express their perspective on a specific Topic to showcase their distinct personality trait.
When we pose questions to LLMs about the Topic: Coldplay using the template “What is your opinion of Coldplay?”, LLMs might respond with vague and inconsistent statements. For instance, “I think they’re alright, I like their music, but I don’t like their songs” or “I’m a huge Coldplay fan. I have to say, I think they’re one of the best bands.” It can be found the first answer depicts an unpredictable sentiment intensity and the model exhibits contradictory viewpoints in the above two responses, which is unsatisfactory. The objective of our proposed task, editing personality for LLMs, aims to modify the model and make it provide responses reflecting a more clear-cut and consistent personality trait. To be specific, if we consider the behaviors of personality trait Neuroticism, an edited model might respond like, “Sometimes the popularity and hype around Coldplay make me feel a little overwhelmed.
2.2 Task Definition
Following model editing  [28, 25, 45], we define the proposed task of editing personality for LLMs as editing the base model $f_{b}$ to the edited model $f_{e}$ with an edit descriptor. Specifically, the basic model $f_{b}$ is represented by a function $f:\mathbb{X}\Rightarrow\mathbb{Y}$ that projects an input $x$ to its corresponding prediction $y$.
Personality Facet Opinion Text
Extraversion assertiveness I believe Arras is worth checking out because of its unique blend of history and culture. You won’t be disappointed with what it has to offer.
@@ 2.2 Task Definition
Table 1: An example of our benchmark PersonalityEdit for the Topic Arras.
@@ 2.3 Comparison with Prior Tasks
Previous model-editing tasks have largely focused on editing factual knowledge within LLMs [28, 27, 25, 24]. This line of work, which includes fact-checking, knowledge editing, and counterfactual model editing, addresses the issue of outdated knowledge within LLMs. The goal of the factual knowledge editing task is to produce an edited model $f_{e}$ such that $f_{e}(x)=y_{e}$ when $x$ is in scope, and $f_{e}(x)=f_{b}(x)$ when $x$ is out-of-scope [45]. Different from factual knowledge editing, our proposed task presents a straightforward editing scope. The edit descriptor in our task is defined by the topic and target personality.
In addition to knowledge editing, [28] proposes the editing of conversational sentiment (ConvSent) within a dialogue agent on a specific topic. Drawing inspiration from this and the research on human personality, we incorporate ConvSent’s focus on responding to specific topics to define the proposed task. However, rather than the binary approach of positive and negative sentiments, we introduce a nuanced personality framework. Our framework features three major personality traits and corresponding facets, facilitating a more granular exploration of personality behavior for LLMs. Besides, text style transfer, such as altering text formality [20, 47] or politeness [22], typically involve transitioning from source text to another while preserving the content. However, our task is proposed on model editing, aiming to gain modified models that can more precisely meet customizing needs about the viewpoints on the specific topic.
Figure 2: Overview of our PersonalityEdit benchmark construction, including selecting personality traits, topic filtering, data generation, and quality control.
@@ 3.1 Selection of Personality Traits and Facets
The field of personality theory encompasses a multitude of studies and definitions of personality [37, 9, 4]. Prominent among these are the Myers-Briggs Type Indicator (MBTI [29]) and the Big Five Personality Traits [9]. The latter, widely recognized for its comprehensiveness, includes Neuroticism, Extraversion, Openness to Experience, Agreeableness, and Conscientiousness.
In conventional discourse or lines from a script, it is feasible to discern multiple dimensions of an individual’s personality traits. For instance, in the previous dataset [15] dedicated to personality recognition, a single text passage typically contains labels across five personality traits. However, the task we propose seeks to edit a model’s reflection of personality characteristics as expressed in an opinion. Thus, our selection of personalities is based on two criteria: 1. The clarity with which personality traits manifest in opinion text; 2. Their distinctiveness from other personality viewpoints, aids in the evaluation of editing outcomes. Note that Extraversion and Neuroticism are the best-understood personality traits in terms of their underlying processes [5], and exhibit more prominent characteristics. They demonstrate clear differentiation from the other three traits. From the remaining, after a detailed analysis, we select Agreeableness,
Items Train Dev Test
#Topics 1,600 200 200
#Average popularity of topics (views) 58107.6 60262.4 56924.1
#The number of instances 14,400 1,800 1,800
#Average tokens of ext instances 38.28 38.65 38.20
#Average tokens of arg instances 43.57 43.90 43.01
#Average tokens of neu instances 43.96 43.78 42.84
#Average tokens of all instances 41.93 42.11 41.35
Table 2: Statisitc for PersonalityEdit benchmark.
as it demonstrated greater distinctiveness in expressing viewpoints compared to the others, to construct our benchmark.
However, the behavior of these traits could result in a simple expression of emotion, similar to previous work in ConvSent [28]. To circumvent this, following [16], we employ the NEO PI-R facets to delineate each personality trait. A facet represents a specific and unique element within a broader personality trait. Facets of Neuroticism include anxiety and depression, while excitement-seeking and gregariousness are facets of Extraversion. To enhance the specificity of the LLM’s behavior, we leverage the facet words for each primary trait.
@@ 3.2 Data Generation
The data construction centers on guiding GPT-4 [30] to generate responses aligned with a specified topic, while also embodying the target personality. The first step is to select suitable topics. Note that previous work [23] indicates that LLMs tend to provide superior responses to topics of high popularity. Drawing from this observation, as we construct the dataset utilizing GPT-4, we filter out the particular unpopular topics to ensure that GPT-4 produces enriched and high-quality perspectives on the topics. We adopt the implementation in [23] to quantify topic popularity and filter out those with low popularity. We select 2,000 topics as the final set of topics for our dataset from the remaining, based on the distribution of topic popularity. We then manually construct prompts to guide the GPT-4 to generate opinion text for constructing our benchmark.
Quality Control. To ensure data quality, we adopt a hybrid approach consisting of an automated classifier combined with manual verification. Specifically, we initially instruct GPT-4 to produce data for 200 topics across three personalities. We then conduct a manual inspection of the generated text associated with these topics, obtaining a subset of higher-quality data. As for the inter-annotator agreement, we provide examples and a list of personality traits, their corresponding facets, and associated adjectives for the annotators. Annotators are asked to assess whether the generated data accurately reflected the designated facet or adjective descriptions and whether there were any ambiguities present. The refined dataset is then used to train a RoBERTa-Base model [21] as the personality classifier. The classifier is subsequently employed for automatic filtering in the following generation and the final evaluation.
“Neuroticsim”: Respond to match the description. Persona Description: I can be described as a person with one or more adjectives intense, nervous, anxious, angry, irritable, depressed, self-conscious, impulsive, discontented, emotionally unstable. Evaluating the opinion: “{}”. how accurately the opinion matches the description, please rate a scale in [1,2,3,4,5] (where 1 = ‘very inaccurate‘, 2 = ‘moderately inaccurate‘, 3 = ‘neither accurate nor inaccurate‘, 4 = ‘moderately accurate‘, and 5 = ‘very accurate‘):
Table 3: Example prompt instructing GPT-4 for evaluation on the generated sentence, consisting of instruction, Persona Description for selected personality, the generated sentence and the statement of evaluation scores.
@@ 4 Experiments Setup
Backbones and Metrics.
@@
ES. and DD. These metrics rely on pre-generated text in the dataset, calculated by the likelihood of the edited model. ES is designed to focus on the inner topic $I(t_{e})$, and DD metric concentrates on the outer topics $O(t_{e})$.
Accuracy. For the opinion text generated $y^{\prime}_{e}$ from edited model $f_{e}$, we employ $p^{\prime}_{e}=\text{{PT}}(y^{\prime}_{e})$ to evaluate the Accuracy on target personality $p_{e}$.
TPEI. We further propose a new metric named Target Personality Edit Index (TPEI). We utilize cross-entropy to measure if $y^{\prime}_{e}$ from $f_{e}$ leans more towards the target personality, compared to the opinion text $y^{\prime}_{b}$ from the base model $f_{b}$, denoted as $cross(.,.)$, and TPEI is formulated as:
PAE. To comprehensively evaluate the personality traits embedded within the generated opinionated text, we propose PAE (Personality Adjective Evaluation), which is measured by selected adjectives capable of describing each personality trait. By modeling our approach after the evaluation questionnaire presented in  [34], we construct prompts for each segment of generated text. GPT-4 assigns a score ranging from 1 to 5 for each generated text segment based on the target personality $p_{e}$, formulated as $\mathrm{pae}\left(y,p_{e}\right)$. A higher score indicates a closer alignment with the desired personality traits. An example prompt is provided in 3 To be specific, the PAE result is calculated by the $y^{\prime}_{e}$ and $y^{\prime}_{b}$ as follows:
@@
Table 4: The main result of the baselines on PersonalityEdit. The $\uparrow$ indicates the metric goes higher if the editing method performs better, and $\downarrow$ indicates the lower the better. We do not report the results of SERAC (The metrics based on generated text are set to ‘-’) because it fails to generate fluent text after editing. We also do not report the MEND result of llama-2-13b-chat as well as both the MEND and SERAC result of llama-2-70b-chat due to the lack of GPU resources. Note that after training with MEND, the edited model cannot always produce fluent text, so we filter out the incoherent cases and report the results.
Baselines.
MEND [27] is a method for implementing local edits to LLMs using a single input-output pair. SERAC [28] provides a technique that channels modified information through a distinct parameter set, thus preserving the initial weights. IKE [50] modifies factual knowledge through In-Context Learning. For the proposed task, IKE is adapted as a straightforward ICL approach. PROMPT, we use a designed prompt to instruct the behaviors of LLMs.
5 Results
5.1 Main Result
From Table 4, it is evident that for the results from the same editing method, the metrics for ES and DD perform better on methods that require training, i.e. MEND and SERAC. This can be attributed to the fact that both MEND and SERAC are optimized based on the loss at the logits level. However, MEND and SERAC challenge to generate fluent text. Specifically, SERAC, by being an external small model add-on, struggles to produce complete sentences. Similarly, when editing llama-2-7b-chat by MEND, it can fail to generate fluent text. Even after filtering out the incoherent cases, it doesn’t exhibit satisfactory editing outcomes
(with an accuracy of merely 29.8% post-editing). When fluent text is generated on GPT-J, the resultant metrics based on generation are not particularly high, suggesting that the ES and DD metrics may not be entirely reliable for assessing opinion text editing tasks. It indicates that this type of metric is not quite reliable as the measure for generating tasks. Additionally, MEND and SERAC do not consistently generate fluent text, especially on aligned models. In contrast to these training-dependent methods, prompt-based editing approaches, i.e., IKE and PROMPT, can generate superior text and achieve better results on generation metrics. This indicates the need for future research for methods that can edit model personality traits or other features without compromising the text generation capabilities of LLMs. Furthermore, it is observed that PROMPT’s editing performance on GPT-J is relatively suboptimal, whereas IKE demonstrates a more consistent performance. On the aligned Llama-2-chat series models, both PROMPT and IKE show markedly better editing success compared to their unaligned counterparts. Besides, the performance gap between PROMPT and IKE narrows as the model’s parameter size increases, aligning with the scaling laws.
Our experiments are confined to the GPT-J and LLaMA2 series models. The results may be different in other LLMs, but our dataset is compatible with other models and alternative editing methods, offering avenues for future work.